# Is Large-scale Pretraining the Secret to Good Domain Generalization?

### v2

@@ Abstract
Multi-Source Domain Generalization (DG) is the task of training on multiple source domains and achieving high classification performance on unseen target domains. Recent methods combine robust features from web-scale pretrained backbones with new features learned from source data, and this has dramatically improved benchmark results. However, it remains unclear if DG finetuning methods are becoming better over time, or if improved benchmark performance is simply an artifact of stronger pre-training. Prior studies have shown that perceptual similarity to pre-training data correlates with zero-shot performance, but we find the effect limited in the DG setting. Instead, we posit that having perceptually similar data in pretraining is not enough; and that it is how well these data were learned that determines performance. This leads us to introduce the Alignment Hypothesis, which states that the final DG performance will be high if and only if alignment of image and class label text embeddings is high. Our experiments confirm the Alignment Hypothesis is true, and we use it as an analysis tool of existing DG methods evaluated on DomainBed datasets by splitting evaluation data into In-pretraining (IP) and Out-of-pretraining (OOP). We show that all evaluated DG methods struggle on DomainBed-OOP, while recent methods excel on DomainBed-IP. Put together, our findings highlight the need for DG methods which can generalize beyond pretraining alignment.
@@ 1 Introduction
Domain Generalization (DG) addresses the challenge of enabling AI models to generalize from known domains to unseen ones, a critical task given the inevitable distribution shifts between training and real-world deployment (Saenko et al., 2010). DG pipelines typically consist of three stages: pretraining a model on a large, general dataset; finetuning the model with one or more source domains; and finally evaluating the model on target domains that are distinct from source domains. More and more, DG methods rely on huge-scale foundation models for initialization. Simultaneously, finetuning has increasingly incorporated regularization to prevent catastrophic forgetting. As a result, it remains unclear whether DG adaptation methods are genuinely improving or if enhanced benchmark performance is simply due to stronger pre-training combined with regularization, or even the presence of target domains within the hundred million-scale pre-training data.
In this work, we examine the reliance of recent DG methods on pre-trained features. We focus on CLIP-based models, as they are used in all state-of-the-art DG methods (Addepalli et al., 2024; Cho et al., 2023; Shu et al., 2023; Mao et al., 2024). While prior studies  (Mayilvahanan et al., 2024) have shown that perceptual similarity to pre-training data explains zero-shot performance—referred to as the Image Similarity Hypothesis—we find this relationship to be limited in the DG setting. Despite evidence of target domains being present in pre-training (Figure 6), perceptual similarity alone does not fully explain accuracy in the DG context. We propose that it is not just the presence of similar data in pre-training that matters, but also how well this data was learned. To this end, we introduce the Alignment Hypothesis, which states that pre-trained alignment between image and class embeddings is still predictive of DG performance even after source finetuning. We find that, performance for low alignment samples can be almost 0, while performance for high alignment samples is close to perfect. These results confirm the Alignment Hypothesis. As illustrated in Figure 1, these findings suggest that current DG methods largely fail to learn new, general features from the source data when the pretraining does not already provide a strong alignment.
@@
The confirmation of the Alignment Hypothesis gives us a tool to separate in-pretraining (IP) data from out-of-pretraining (OOP) data for a particular backbone, and we do so for five DG datasets with OpenCLIP-ViT/B-16. We call the resulting splits DomainBed-IP and DomainBed-OOP. Evaluating on DomainBed-IP/OOP offers a view of where current DG methods fail and where they succeed. We find that all methods, including those considered state-of-the-art, perform poorly on OOP data, i.e. data that the pretrained backbone hadn’t already aligned well. Furthermore, recent state-of-the-art methods do not outperform older methods on OOP data. For example, CLIPood  (Shu et al., 2023) slightly under performs a combination of older methods (MIRO (Cha et al., 2022) + MPA (Arpit et al., 2022)) on DomainBed-OOP. At the same time, existing DG methods show exceptional performance on DomainBed-IP, sometimes even outperforming an oracle model trained on the target domain. These results suggest that future research should aim to enhance DG methods on low-alignment data while preserving the already strong performance on high-alignment data. In summary, we make the following contributions:
@@
  • Propose a new IP/OOP evaluation framework: We demonstrate that splitting target data by its alignment with the pre-trained backbone can effectively test Out-of-Pretraining (OOP) generalization. We will release IP/OOP splits for the DomainBed dataset to support future research.
@@ 2 Related Work
Multi-Source Domain Generalization: Domain Generalization aims to mitigate the impacts of domain shifts between source (training) and target (deployment) domains. One standard approach is domain-invariant feature learning, which leverages domain labels to learn domain-invariant features. CORAL  (Sun & Saenko, 2016) aligns second-order statistics, while DANN  (Ganin et al., 2016) uses an adversarial loss.  Gulrajani & Lopez-Paz (2020) show that ERM, which does not align features between domains, can outperform most prior work while being easier to tune. Another common approach is domain-aware data augmentation to expand the training domain to become closer to or even overlap the target domain. Inter-domain mixup  (Yan et al., 2020) blends images from different domains. Similarly, style transfer can diversify training images  (Zhong et al., 2022). Deep ensembles are effective for domain generalization (Arpit et al., 2022). Since they are computationally inefficient for inference, many recent works average model weights from either multiple finetuning runs or from a single training trajectory (Cha et al., 2021; Arpit et al., 2022; Rame et al., 2022; Jain et al., 2023; Li et al., 2023; Shu et al., 2023). More recently, several methods perform regularized finetuning towards the initialization of a pretrained model. This works under the assumption that pretrained features are useful for target data, and should not be unlearned. The general idea can be applied to weight space (L2SP (Xuhong et al., 2018)), feature space (MIRO (Cha et al., 2022)), or output space (CAR-FT (Mao et al., 2024),CLIPood (Shu et al., 2023)).
@@
Impact of Data on Model Performance: Several recent studies have explored the influence of pretraining data on model performance. Mayilvahanan et al. (2024) investigated how the presence of perceptually similar images in CLIP (Radford et al., 2021) pretraining affects performance, introducing the Similarity Hypothesis, which posits that nearest neighbor similarity is strongly correlated with zero-shot accuracy. Udandarao et al. (2024) demonstrated that concept frequency in pretraining is correlated with zero-shot performance and introduced a dataset focusing on infrequent concepts. Fang et al. (2022) found that diversity in pretraining data is critical for improving performance on benchmarks such as ImageNetV2 (Recht et al., 2019), ImageNet-R (Hendrycks et al., 2021), ImageNet-Sketch (Wang et al., 2019), and ObjectNet (Barbu et al., 2019). However, these studies focus on the zero-shot setting, where models are evaluated without further training. In contrast, we examine the domain generalization setting, where pre-trained models are fine-tuned on source domains and tested on held-out target domains. Our findings suggest that comparing target images to pre-trained images, as proposed by Mayilvahanan et al. (2024), is less predictive of final DG performance than directly measuring the alignment between the image and its class embedding.
@@ 3.1 Testing the Image Similarity Hypothesis
Figure 2: Comparing the Predictive Power of the Alignment and Image Similarity Hypotheses for Domain Generalization (DG). a.) Alignment Hypothesis: The cosine similarity between image and ground truth text-label embeddings before pre-training (Alignment Score) is highly predictive of model accuracy after fine-tuning on five DG datasets, with Alignment Score distributions shown in the colored histograms. This suggests that image-text pairs well-aligned during pre-training result in better performance on target tasks. b.) Image Similarity Hypothesis: In contrast, the cosine similarity between a test image and its closest match from the pre-training set (Perceptual Similarity Score) shows weaker predictive power for accuracy, implying that visual resemblance alone is less indicative of downstream performance.
@@ 3.2 Introducing the Alignment Hypothesis
$\text{AlignmentScore}(I,T)=\frac{\langle f_{I}(I),f_{T}(T)\rangle}{\|f_{I}(I)\|\cdot\|f_{T}(T)\|}$
where $f_{I}(I)$ is the embedding of the image before finetuning on source, and $f_{T}(T)$ is the embedding of the text.
We verify the Alignment Hypothesis similarly to the Image Similarity Hypothesis, by binning samples using the AlignmentScore and computing accuracy for each bin using VL2V-SD. We provide the same analysis for many more DG methods in the Appendix Figure 8. In Figure 2 a.), we can see that the Alignment Hypothesis explains DG performance after source finetuning, significantly more strongly than for the Image Similarity Hypothesis in Figure 2 b.) This finding suggests that source fine-tuning in DG, which aims to achieve high performance across all target samples, only succeeds on those with high initial alignment.
@@ 4 Re-thinking Domain Generalization Benchmarking using the Alignment Hypothesis
Knowing that the Alignment Hypothesis holds for contrastively trained image-text models (Section 3), we can now use it as a tool to probe the performance of DG methods across different levels of pre-training alignment. We apply this approach to five widely-used DomainBed (Gulrajani & Lopez-Paz, 2020) DG datasets: VLCS (Fang et al., 2013), PACS (Li et al., 2017), OfficeHome (Ganin et al., 2016), TerraIncognita (Beery et al., 2018), and DomainNet  (Peng et al., 2019). This section discusses how we create new splits for existing DG datasets using AlignmentScore. We start by computing AlignmentScore for all samples in 5 DG datasets (Figure 3). Based on our observation that some samples are mislabeled, we perform dataset cleaning (4.1). Finally, we find an AlignmentScore threshold to split DG datasets into In-Pretraining (IP) and Out-of-Pretraining (OOP) evaluation subsets (Section 4.2), which we later use for evaluating DG methods (Section 5). In order to connect the AlignmentScore with the DG method, we use the same backbone both for splitting the datasets into IP and OOP subsets and for training DG methods.
@@ 4.1 Data Exploration and Cleaning
We start by visualizing the data of all the datasets at various AlignmentScore values. We show some representative samples in Figure 3. At very low scores, we find that a large fraction of labels are incorrect (red boxes in Figure 3).
As a result, we divide the data into AlignmentScore intervals (e.g., 0.00-0.05, 0.05-0.10, and so on, up to 0.2) and randomly sample 100 instances from each interval for every dataset. This allows us to systematically analyze the relationship between AlignmentScore and label accuracy across different score ranges. For each interval, we then count the fraction of mislabelled samples to better understand how low AlignmentScores are associated with labeling errors. We find that below an AlignmentScore of 0.15, label noise is unacceptably high, with all datasets suffering the most from mislabelling (Table 4 in Appendix). Therefore, we discard all samples with AlignmentScore less than 0.15 in DomainBed-IP/OOP. As shown in Table 1, we observe that the percentage of discarded samples due to mislabeling varies across datasets, with VLCS and DomainNet having the highest rates at 12.41% and 7.64%, respectively.
Furthermore, on the right side of Figure 3, we observe that at very high AlignmentScores (greater than 0.4), the images often contain text directly related to the label. Since our goal is to evaluate visual recognition rather than text recognition (OCR), and CLIP is known to have strong OCR abilities  (Fort, 2021), we exclude all samples above AlignmentScore of 0.4 from DomainBed-IP/OOP. As shown in Table 1, although only a small portion of data is removed due to OCR filtering (0.00-0.15% across datasets), this issue may become more significant in future studies.
Dataset Dropped - Mislabelled Dropped - OCR
OfficeHome 0.92% 0.15%
PACS 3.05% 0.00%
TerraIncognita 0.22% 0.00%
DomainNet 7.64% 0.03%
VLCS 12.41% 0.00%
@@ 4.2 Data Splitting
After filtering, we focus on determining a threshold to split the dataset into In-Pretraining (IP) and Out-of-Pretraining (OOP) subsets. We select 0.21 as the threshold, based on the trends observed in Figure 2 (a.)), as this is the point where performance begins to improve significantly, indicating that existing methods become more effective. While this threshold represents a somewhat subjective choice informed by observed patterns, we provide AlignmentScores in the released data, allowing researchers the flexibility to experiment with their own thresholds.
Figure 4 illustrates how this split impacts the size and composition of each dataset. Certain domains naturally fall into the IP subset. For example, in DomainNet, the clipart domain is predominantly categorized as IP, likely due to its frequent presence on the internet and therefore the web-scraped pre-training data. On the other hand, TerraIncognita-OOP has relatively more balanced domains, but exhibits substantial class shift between OP and OOP splits (Figure 5), meaning some classes are better aligned than others during pre-training. More dataset statistics are in the Appendix.

### v3

@@ Abstract
Multi-Source Domain Generalization (DG) is the task of training on multiple source domains and achieving high classification performance on unseen target domains. Recent methods combine robust features from web-scale pretrained backbones with new features learned from source data, and this has dramatically improved benchmark results. However, it remains unclear if DG finetuning methods are becoming better over time, or if improved benchmark performance is simply an artifact of stronger pre-training. Prior studies have shown that perceptual similarity to pre-training data correlates with zero-shot performance, but we find the effect limited in the DG setting. Instead, we posit that having perceptually similar data in pretraining is not enough; and that it is how well these data were learned that determines performance. This leads us to introduce the Alignment Hypothesis, which states that the final DG performance will be high if and only if alignment of image and class label text embeddings is high. Our experiments confirm the Alignment Hypothesis is true, and we use it as an analysis tool of existing DG methods evaluated on DomainBed datasets by splitting evaluation data into In-pretraining (IP) and Out-of-pretraining (OOP). We show that all evaluated DG methods struggle on DomainBed-OOP, while recent methods excel on DomainBed-IP. Put together, our findings highlight the need for DG methods which can generalize beyond pretraining alignment. We release DomainBed-OOP at https://huggingface.co/datasets/PTeterwak/DomainBed_OOP.
@@ 1 Introduction
Domain Generalization (DG) addresses the challenge of enabling AI models to generalize from known domains to unseen ones, a critical task given the inevitable distribution shifts between training and real-world deployment (Saenko et al., 2010). DG pipelines typically consist of three stages: pretraining a model on a large, general dataset; finetuning the model with one or more source domains; and finally evaluating the model on target domains that are distinct from source domains. DG methods increasingly rely on huge-scale foundation models for initialization (e.g.(Shu et al., 2023; Cha et al., 2022; Addepalli et al., 2024)). Simultaneously, finetuning has increasingly incorporated regularization to prevent catastrophic forgetting. As a result, it remains unclear whether DG methods are genuinely improving or if benchmark performance gains are simply due to stronger pre-training, possibly with the target domains within the hundred million-scale pre-training data (Mayilvahanan et al., 2025), combined with regularization.
In this work, we examine the reliance of recent state-of-the-art CLIP-based DG methods on pre-trained features. While prior studies have shown that perceptual similarity to pre-training data explains zero-shot performance—referred to as the Image Similarity Hypothesis (Mayilvahanan et al., 2024)—we find this relationship to be limited in the DG setting. Despite evidence of target domains being present in pre-training (Figure 4), we find that perceptual similarity alone does not fully explain accuracy in the DG context (Section 3). We argue that it is not just the presence of similar data in pre-training that matters, but also how well this data was learned. To this end, we introduce the Alignment Hypothesis, which states that pre-trained alignment between image and class embeddings is still predictive of DG performance even after source finetuning. We note that we do not make assumptions of how or why alignment arose. In the Alignment Hypothesis, pre-training alignmnent is used as a measure for how well a sample is learned. We find that performance for low alignment samples can be almost 0, while performance for high alignment samples is close to perfect. These results confirm the Alignment Hypothesis. As illustrated in Figure 1, these findings suggest that current DG methods largely fail to learn new, general features from the source data when the pretraining does not already provide a strong alignment.
@@
The confirmation of the Alignment Hypothesis gives us a tool to separate aligned and well learned in-pretraining (IP) data from misaligned and poorly learned out-of-pretraining (OOP) data for a particular backbone, and we do so for five DG datasets with OpenCLIP-ViT/B-16. We call the resulting splits DomainBed-IP and DomainBed-OOP. Evaluating on DomainBed-IP/OOP offers a view of where current DG methods fail and where they succeed. We focus on CLIP-based methods, as they are used in state-of-the-art DG methods (Addepalli et al., 2024; Cho et al., 2023; Shu et al., 2023; Mao et al., 2024); we believe extensions to pure vision models such as DINOv2 (Oquab et al., 2023) represent interesting future work. We find that all methods, including those considered state-of-the-art, perform poorly on OOP data, i.e., data that the pretrained backbone hadn’t already aligned well. Furthermore, recent state-of-the-art methods do not outperform older methods on OOP data; CLIPood  (Shu et al., 2023) slightly under performs a combination of older methods (MIRO (Cha et al., 2022) + MPA (Arpit et al., 2022)) on DomainBed-OOP. At the same time, existing DG methods show exceptional performance on DomainBed-IP, sometimes even outperforming an oracle model trained on the target domain. These results suggest that future research should aim to enhance DG methods on low-alignment data. In summary, we make the following contributions:
@@
  • Propose a new IP/OOP evaluation framework: We demonstrate that splitting target data by its alignment with the pre-trained backbone can effectively test Out-of-Pretraining (OOP) generalization. We release IP/OOP splits for DomainBed datasets.
@@ 2 Related Work
Multi-Source Domain Generalization: Domain Generalization aims to mitigate the impacts of domain shifts between source (training) and target (deployment) domains. These can include subpopulation shifts where all testing subpopulations are present in the training data but in different proportions (Dehdashtian et al., 2024), or it could be the case we consider in this work where the testing subpopulation is not at all present in the training subpopulation. Although we focus on the multi-source domain generalization task, where domains between train and test are completely disjoint, our analysis can be extended to other types of generalization. One standard approach is domain-invariant feature learning, which leverages domain labels to learn domain-invariant features. CORAL (Sun & Saenko, 2016) aligns second-order statistics, while DANN (Ganin et al., 2016) and ADDA (Tzeng et al., 2017) uses an adversarial loss.  Gulrajani & Lopez-Paz (2020) show that ERM, which does not align features between domains, can outperform most prior work while being easier to tune. Another common approach is domain-aware data augmentation to expand the training domain to become closer to or even overlap the target domain. Inter-domain mixup (Yan et al., 2020) blends images from different domains. Similarly, style transfer can diversify training images (Zhong et al., 2022). Deep ensembles are effective for domain generalization (Arpit et al., 2022). Since they are computationally inefficient for inference, many recent works average model weights from either multiple finetuning runs or from a single training trajectory (Cha et al., 2021; Arpit et al., 2022; Rame et al., 2022; Jain et al., 2023; Li et al., 2023; Shu et al., 2023). More recently, several methods perform regularized finetuning towards the initialization of a pretrained model. This works under the assumption that pretrained features are useful for target data, and should not be unlearned. The general idea can be applied to weight space (L2SP (Xuhong et al., 2018)), feature space (MIRO (Cha et al., 2022)), or output space (e.g., CAR-FT (Mao et al., 2024), CLIPood (Shu et al., 2023)).
@@
Impact of Data on Model Performance: Several recent studies have explored the influence of pretraining data on model performance.  Mayilvahanan et al. (2024) investigated how the presence of perceptually similar images in CLIP (Radford et al., 2021) pretraining affects performance, introducing the Similarity Hypothesis, which posits that nearest neighbor similarity is strongly correlated with zero-shot accuracy.  Mayilvahanan et al. (2025) show that domain contamination in the pre-training has substantial impact on DG performance.  Udandarao et al. (2024) demonstrated that concept frequency in pretraining is correlated with zero-shot performance and introduced a dataset focusing on infrequent concepts.  Fang et al. (2022) found that diversity in pretraining data is critical for improving performance on benchmarks such as ImageNetV2 (Recht et al., 2019), ImageNet-R (Hendrycks et al., 2021), ImageNet-Sketch (Wang et al., 2019), and ObjectNet (Barbu et al., 2019). However, these studies focus on the zero-shot setting, where models are evaluated without further training. In contrast, we examine the domain generalization setting, where pre-trained models are fine-tuned on source domains and tested on held-out target domains.  Yu et al. (2024) recommend using self-supervised pre-training to avoid data label leakage. In contrast we study DG model behavior in the more realistic setting of CLIP-pretraining. Our findings suggest that comparing target images to pre-trained images, as proposed by  Mayilvahanan et al. (2024), is less predictive of final DG performance than directly measuring the alignment between the image and its class embedding.
@@ 3.1 Testing the Image Similarity Hypothesis
Figure 2: Comparing the Predictive Power of the Alignment and Image Similarity Hypotheses for Domain Generalization (DG). a.) Image Similarity Hypothesis: The cosine similarity between a test image and its closest match from the pre-training set (Perceptual Similarity Score) shows relatively weak predictive power for accuracy, implying that visual resemblance alone is not fully indicative of downstream performance. b.) Alignment Hypothesis: In contrast, the cosine similarity between image and ground truth text-label embedding after pre-training (Alignment Score) is highly predictive of model accuracy after fine-tuning on five DG datasets, with Alignment Score distributions shown in the colored histograms. This suggests that image-text pairs well-aligned during pre-training result in better performance on target tasks.
@@
Figure 2 a.) shows the results of this analysis of the recent, high-performing VL2V-SD (Addepalli et al., 2024) on various DG datasets. While the Image Similarity Hypothesis is somewhat predictive of DG performance, its influence is not very strong. This suggests that perceptually similar pretraining data alone may not guarantee high DG performance; additional factors, such as how effectively the nearest neighbors were aligned with the target concept, may also be significant.
Algorithm 1 Evaluating the Image Similarity Hypothesis
1:Target domain samples $D_{\text{target}}$, trained DG model $M$, pre-trained image encoder $f_{I}$,
2:for each sample $I\in D_{\text{target}}$ do
3:     Retrieve nearest neighbor of $I$ in LAION-400M using $f_{I}$ features, assign to $I_{k}$
4:     Compute PerceptualSimilarityScore($I,I_{k}$) using Equation Eq. 1,
5:     Record correctness of $M(I)$
6:end for
7:Bin samples based on PerceptualSimilarityScore
8:Compute DG accuracy within each bin
9:return Accuracy for each bin
@@ 3.2 Introducing the Alignment Hypothesis
$\text{AlignmentScore}(I,T)=\frac{\langle f_{I}(I),f_{T}(T)\rangle}{\|f_{I}(I)\|\cdot\|f_{T}(T)\|}$ (Eq. 2)
where $f_{I}(I)$ is the embedding of the image before finetuning on source, and $f_{T}(T)$ is the embedding of the text. Relative to directly using the contrastive loss value which also depends on negatives, PerceptualSimilarityScore has a scale which aligns across datasets (Appendix B.6).
We verify the Alignment Hypothesis similarly to the Image Similarity Hypothesis, by binning samples using the AlignmentScore and computing accuracy for each bin using VL2V-SD. We provide the same analysis for many more DG methods in the Appendix Figure 11. In Figure 2 b.), we can see that the Alignment Hypothesis explains DG performance after source finetuning, significantly more strongly than for the Image Similarity Hypothesis in Figure 2 a.) This finding suggests that source fine-tuning in DG, which aims to achieve high performance across all target samples, only succeeds on those with high initial alignment.
@@ 4 Re-thinking Domain Generalization Benchmarking using the Alignment Hypothesis
Knowing that the Alignment Hypothesis holds for contrastively trained image-text models (Section 3), we can now use it as a tool to probe the performance of DG methods across different levels of pre-training alignment. We apply this approach to five widely-used DomainBed (Gulrajani & Lopez-Paz, 2020) DG datasets: VLCS (Fang et al., 2013), PACS (Li et al., 2017), OfficeHome (Ganin et al., 2016), TerraIncognita (Beery et al., 2018), and DomainNet  (Peng et al., 2019). This section discusses how we create new splits for existing DG datasets using AlignmentScore. We start by computing AlignmentScore for all samples in 5 DG datasets (Figure 3). Based on our observation that some samples are mislabeled, we perform dataset cleaning (Section 4.1). Then we find an AlignmentScore threshold to split DG datasets into well aligned In-Pretraining (IP) and poorly aligned Out-of-Pretraining (OOP) evaluation subsets (Section 4.2), which we later use for evaluating DG methods (Section 5). In order to connect the AlignmentScore with the DG method, we use the same backbone both for splitting the datasets into IP and OOP subsets and for training DG methods.
@@ 4.1 Data Exploration and Cleaning
We start by visualizing the data of all the datasets at various AlignmentScore values. We show some representative samples in Figure 3. At very low scores, we find that a large fraction of labels are incorrect (red boxes in Figure 3). Thus, we divide the data into AlignmentScore intervals (e.g., 0.00-0.05, 0.05-0.10, and so on, up to 0.2) and randomly sample 100 instances from each interval for every dataset. This allows us to systematically analyze the relationship between AlignmentScore and label accuracy across different score ranges. For each interval, we then count the fraction of mislabelled samples to better understand how low AlignmentScores are associated with labeling errors. We find that below an AlignmentScore of 0.15, label noise is unacceptably high, with all datasets suffering the most from mislabelling (Table 7 in Appendix). Therefore, we discard all samples with AlignmentScore less than 0.15 in DomainBed-IP/OOP. As shown in Table 8 (in the Appendix), we observe that the percentage of discarded samples due to mislabeling varies across datasets, with VLCS and DomainNet having the highest rates at 12.41% and 7.64%, respectively.
Furthermore, on the right side of Figure 3, we observe that at very high AlignmentScores (greater than 0.4), images often contain text directly related to the label. However, our goal is to evaluate visual recognition rather than text recognition (OCR), and CLIP is known to have strong OCR abilities (Fort, 2021), so we exclude all samples above AlignmentScore of 0.4 from DomainBed-IP/OOP. As shown in Table 8 (in the Appendix), only a small portion of data is removed due to OCR filtering (0.00-0.15% across datasets), but this issue may become more significant in future studies.
@@ 4.2 Data Splitting
After filtering, we focus on determining a threshold to split the dataset into In-Pretraining (IP) and Out-of-Pretraining (OOP) subsets. We select 0.21 as the threshold, based on the trends observed in Figure 2 b.), as this is the point where performance begins to improve significantly, indicating that existing methods become more effective. While this threshold represents a somewhat subjective choice informed by observed patterns, we provide AlignmentScores in the released data, allowing researchers the flexibility to experiment with their own thresholds.
Figure 18 in the Appendix shows how this split impacts the size and composition of each dataset. For example, clipart in DomainNet is predominantly categorized as IP, likely due to its frequent presence on the internet and, therefore, web-scraped pre-training data. In comparison, TerraIncognita-OOP is more balanced domains, but exhibits substantial class shift between IP and OOP splits (Figure 17 in the Appendix), meaning some classes are better aligned than others during pre-training.
We further investigate the distinctions between the IP and OOP subsets using VisDiff (Dunlap et al., 2024), an LLM-based system that identifies differences between sets of images. For each combination of dataset, domain, and class—except for DomainNet, where we subsample roughly 15% of the combinations due to computational constraints—we independently sample up to 30 images from both the IP and OOP subsets. VisDiff employs CLIP (Radford et al., 2021) to compute an AUC-ROC score for the natural language differences proposed by the LLM, and we retain only those differences with an AUC score of 0.7 or higher. Results are presented in Appendix Table LABEL:tab:visdiff.
Several clear patterns emerge from our analysis. Contextual or environmental elements often overshadow the primary object indicated by the text label. For example, in VLCS’s SUN09 domain, the OOP subset for the car class frequently features images described as “historical architecture,” while in Office Home the real-domain bed images are deemed OOP if they include “child-themed decor.” These findings imply that object size may also play a role. In the car example, visible architecture implies that the car occupies only a small portion of the image, and in the bed example, the presence of child-themed decor indicates that the entire bedroom is visible rather than just the bed. To test this hypothesis, we computed the object size of the ground truth labeled objects using bounding boxes generated by the open-world object detector OWLv2 (Minderer et al., 2023). We found object size is correlated with the AlignmentScore (see Appendix Figure 10), with the sole exception being TerraIncognita. This indicates that even the presence of small wildlife can substantially increase the AlignmentScore, and that the background vegetation does not present a significantly conflicting signal that is sometimes present in other datasets. Overall, our findings suggest that pre-training which better represents scenes with multiple objects or concepts may improve benchmark performance.